OpenStudy (anonymous):

can someone explain this to me. verify that the numbers 1+sqrt 5 and 1- sqrt 5 both satisfy the equation x^2-2x-4=0

OpenStudy (anonymous):

let x = 1+sqrt(5) and see what happens. for example, (1+sqrt(5))^2 = (1+sqrt(5))(1+sqrt(5)) = 1+ 2sqrt(5) + 5 = 6+2sqrt(5) by distributing

OpenStudy (ash2326):

we have x^2-2x-4=0 let's find the roots of this one standard quadratic equation ax^2+bx+c=0 \[x=(-b\pm\sqrt{b^2-4ac})/2a\] here a =1 b=-2 c=-4 \[x=(2\pm\sqrt{4+16})/2\] or \[x=1\pm\sqrt 5\] hence the two no. satisfy the ewuation

OpenStudy (ash2326):

*equation